# Can one size fit all?: Measuring Failure in Multi-Document Summarization Domain Transfer

### v1

@@ 4 Results
We analyze the results in the framework of our three axes: dataset domain, model type, and evaluation metric. For ease of analysis, we show the relative performance of each model when evaluated on a new domain in Tables 2 and 1. The relative performance is the percentage gain or loss of a model’s performance when compared to the performance on the in-domain data. The zero-shot baseline with Llama 3.1 is compared to the in-context Llama 3.1 performance for each dataset (i.e., no examples instead of one example). See Appendix C for the detailed per-model and dataset results.
4.1 Domain Generalizability
When aggregated by dataset over the different model types, clear patterns emerge from Table 1.
Science summaries have the lowest general quality.
The quality is lowest in the Science domain compared to Conversation and News according to Fluency and Coherence metrics, even before evaluating Science-trained MDS models on other domains, $0.78$ and $0.64$, respectively. These scores are $10+$ points lower than those of News ($0.91$/$0.79$) and Conversation ($0.91$/$0.72$).
Fluency is affected the least by domain transfer.
The fluency scores are the most stable, i.e., have the least change in performance (Avg $\Delta$) compared to the other metrics. While on average there is a slight decline in fluency, as indicated by the negative change, the decline is negligible in the $(-0.02,-0.01)\%$ range. Fluency does not appear to be an issue with these models, mostly likely due their size (BART-Large and Llama 8B).
Factuality Ground-Truth Similarity Quality
Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm 0.65 0.62 0.32 0.18 0.08 0.83 0.57
Avg. $\Delta$2% -11% -28% -26% -152% +3% -4%
PRIMERA 0.78 0.74 0.35 0.20 0.12 0.89 0.78
Avg. $\Delta$ +3% -10% -19% -24% -16% -14% -1%
SummN 0.66 0.60 0.33 0.19 0.08 0.81 0.56
Avg. $\Delta$ +4% -9% -32% -33% -31% +7% +18%
Llama 3.1 0.82 0.91 0.29 0.18 0.11 0.94 0.94
Avg. $\Delta$ +1% -1% +7% +4% +25% -0.00% +0.00
Llama 3.1 (0-shot) +1% -4% -5% +0% -7% -1% -1%
Table 2: Multi-document summarization (MDS) model performance and relative performance aggregated by model when trained and tested on the same domain (top) and different domains ($\Delta$), respectively. Results are aggregated across all the test splits of each dataset. A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data. A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range. LLama zero-shot is compared to Llama with in-domain examples. Detailed results are in Tables 5 and 6.
There is no clear general-purpose domain for transferring.
Each domain has different benefits and cons to train on for transferring, as indicated by the varied relative performance scores ($\Delta$). Transferring from Conversation to News and Science can lead to greatly increased factuality, as measured by UniEval-Consistency ($+18\%$), but also a decline in Ground-Truth Similarity ($-31\%$ to $-9\%$). Transferring from News appears to be the worst according to the similarity metrics, over $-92\%$ according to MoverScore. The Science-trained models appear to be the most neutral overall, ranging from $-12\%$ to $+6\%$ across all metrics, but claiming that it is the best domain for general MDS model training is difficult because the model scores started lower than the other domains to begin with.
News-trained models have the worst similarity when transferring to non-News data.
There is a steep decline in performance when News-trained models are evaluated on Conversation and Science as indicated by ROUGE ($-51\%$ R1, $-41\%$ RL-Sum) and MoverScore ($-92\%$). However, the Relevance score is high ($+2\%$). The discrepancy in word-mapping metrics (lexical with ROUGE and semantic with MoverScore) and overall similarity (Relevance) show that a summary can still incorporate information from the ground-truth without having an overlap in words.
Conversation and News data have the most unique styles.
The ConvoSumm-Reddit example from Figure 1 shows the stark difference in style between the types of summaries in ConvoSumm-Reddit and those from Multi-News. This difference in style explains the performance loss from Conversation and News domains, as the summary wording would be completely different and penalized by the ground-truth similarity metrics ($-31$ to $-9\%$ and $-92$ to $+2\%$, respectively).
4.2 MDS Model Transfer
Grouping the domain transfer results by model type across all dataset domains reveals even more performance trends (Table 2).
@@ 4.2 MDS Model Transfer
Pre-trained LLM outperforms all fine-tuned smaller models.
For all metrics besides similarity, Llama 3.1 outperforms the other models concerning factuality and quality. This trend holds during domain transfer, with a performance difference ranging from $-1$ to $+25\%$ across all metrics. The second-best performing model across all metrics, including accuracy, is PRIMERA, indicating an advantage of directly fine-tuning a long-context model (4K context window).
@@
Surprisingly, Llama has a similar performance at generating summaries with (in-context learning, ICL) or without being shown an example (zero-shot). This indicates a strong bias to its pre-training versus “learning” from the example. These results could differ with the addition of more examples (at the expense of a smaller context window available for the source documents) or a larger model.
ConvoSumm claim extraction is ineffective on Science.
The first step in the extract-then-summarize model ConvoSumm is to filter out sentences that are not identified as claims by a small, fine-tuned model. This step removes $35\%$ of sentences on average from each ConvoSumm-Reddit example, $51\%$ from Multi-News, but less than $1\%$ from Multi-XScience. This is intuitive since there would be more filler sentences in conversations and news articles, but less so in scientific publications.
@@
With all models, including Llama 3.1 with a 131K context window, there were issues fitting all source documents for every example comfortably in the context windows. As mentioned in Section 3, PRIMERA “handles” this issue by evenly dividing the context window among the source documents and truncating the ends. The assumption that the beginning of documents is more important than the end does not hold for all domains and has only been identified in News (Dey et al., 2020). While ConvoSumm was able to fit all source documents from ConvoSumm-Reddit, it did have to truncate documents from 25 samples in Multi-XScience and 699 samples from MultiNews. PRIMERA, with a context window double the size of ConvoSumm (2K vs 4K), was able to fit all Conversation and Science samples, but truncated 835 Multi-News examples. Llama only had issues with 3 samples from Multi-News.
@@ 4.3 Evaluation of Evaluation
Word-based similarity scores are low across all models.
For all models, the generated summaries are not very similar to the ground-truth reference according to the lexical and semantic similarity metrics ROUGE and MoverScore. The model with the highest accuracy ($0.20$ ROUGE_L and $0.12$ MoverScore) was PRIMERA, higher than the much larger LM, Llama-8B. For lexical-based metrics like ROUGE, the low similarity can be explained through the model’s use of synonyms, which would not be captured through token-based recall. However, the low MoverScores can be explained by (1) the penalization of generations longer or more detailed than the reference and (2) the lack of a normalizing constant as provided by BERTScore (Zhang et al., 2020b). [^5] Zhang et al. (2020b) use a learned constant to normalize the scores to $[0,1]$ for easier readability. Still, low scores across both accuracy metrics indicate that the generated summaries differ from their human-written counterparts. Despite overall low accuracy scores, all models have the highest ROUGE/MoverScore performance on Multi-News (News), indicating the summaries might not be as abstract as the other datasets.
Surprisingly, the Relevance score which compares the reference and generation as a whole instead of token mappings, shows higher similarity, especially for Llama ($0.91$ average).
Accuracy is affected the most.
The reference-based metric, accuracy, shows the most extreme gains and losses compared to the other metrics. This is intuitive because a model’s summary can still be good but not be written in the same manner, or style, as the reference, which can happen if a model is biased towards the writing style of its training data. The zero-shot baseline, which is only prompted and not given any styles, shows the least variance in accuracy compared to the other models.
All models are fluent.
Unsurprisingly, all models have high fluency scores ($0.80+$) since all models are of size BART-Large and above. While fluency is still an important attribute to check, it might not be as important as it once was.
@@
Evaluating for attributes that depend on the source documents (Consistency and Coherence) is most affected by the long-document nature of multi-document summarization. The options are either to 1) truncate the documents, which can greatly alter the scores depending on which documents are included, or 2) calculate the metric for every source document and aggregate the scores. More work is needed to determine whether truncating the source documents for UniEval over- or under-estimates model performance. Another option is to turn to large-context models like GPT-4 to be a “judge” for specific criteria (Kim et al., 2024; Vu et al., 2024). This use of LLMs, which typically can fit 10K+ tokens in their context windows, is a potential solution to this scaling issue. The reference-based metrics, i.e., similarity metrics that only compare the ground-truth and generated summaries, are not affected by this issue.
Trained metrics can penalize out-of-training data.
Just as summarization models can have lower performance on out-of-domain data, so can trained evaluation models like UniEval. As noted before, the fluency and coherence scores on Science are $0.10$ points lower than on News and Conversation. This could be an issue with the summarization model training, but the fluency and coherence scores are also lower on the ground-truth summaries, $0.30+$ points lower (“Ref.” in Table 1), which should not be the case as all summaries are human-written and should have high scores. An explanation is that UniEval was trained on a synthetically augmented version of news data, CNN/DailyMail, which is very different from Scientific text.
@@ 5 Discussion and Conclusion
In most work that surveys or proposes multi-document summarization (MDS) systems, there is little overlap between the analysis of model performance and the shortcomings of popularly used summarization evaluation metrics. We bridged this gap through experiments and analysis of not only the zero-shot domain transfer ability of MDS systems but also their evaluation. We analyzed three components of a multi-document summarization (MDS) pipeline for domain transfer: 1) the domain of the training data, 2) the model type, and 3) the common evaluation metrics.
We found important facets to consider in each component. For domain transfer, the training and test domains are important for model performance, regardless of the model type. While we did not find a clear trend between Conversation, News, and Science, it was apparent that the transfer was highly affected by the style of the summaries, albeit more than the domain itself. Summaries from different domains but written in the same style (e.g., high-level vs low-level) could transfer better with higher similarity to the reference.
For different approaches to MDS, a model that didn’t hold a strong bias towards the style of its training (or example) data is ideal, as is the case with a large pre-trained decoder model, like Llama 3.1. If the domain styles differ greatly, a zero- or few-shot approach with less bias is ideal.
Notable issues we uncovered with evaluation metrics that prevented the fair evaluation of summaries were domain-specific penalizations, mostly for Science, questions with scaling to the MDS case, and the inability of ground-truth similarity metrics to adjust for style change (ROUGE and MoverScore). When evaluating an MDS model on a new domain, word-similarity metrics judge style more than content similarity and trained evaluators could over- or under-judge model performance due to limited information in the context window.
@@ Appendix A Dataset Details
We include three MDS datasets: Multi-News (Fabbri et al., 2019), Multi-XScience (Lu et al., 2020)), and ConvoSumm-Reddit (Fabbri et al., 2021a). We keep the original train, validation, and test splits as the original papers for all datasets. See Table 3 for an example from each dataset and Table 4 for summary statistics.
Multi-News.
For Multi-News we used the “cleaned” version provided by the authors [^6] https://github.com/Alex-Fabbri/Multi-News/issues/36 with 44964/5621/5621 examples for train, validation, and test sets, respectively. We keep the NEWLINE_CHAR symbol in the documents.
@@ Appendix B Training Details
Unless specified otherwise, each model was trained on a single Nvidia A100 GPU with 80GB of memory. The training hyperparameters for PRIMERA, SummN, and ConvoSumm-arg-filtered are in Tables 10, 9 and 11, respectively. Below are the prompts and details for running Llama for the zero-shot and in-context 1-shot comparisons.
@@
Due to the large size of Llama 8B, we hosted it as a service with vLLM on 4 Nvidia A100 80GB GPUs (Kwon et al., 2023). [^7] https://docs.vllm.ai We followed model examples on how to interact for the summarization task (Table 12).
Appendix C Raw Results
@@
The domain-transfer results in Section 4.2 show the summarized relative performance of the models. The detailed performance across domains and models are in Tables 7 and 6, respectively. The raw, un-normalized numbers are shown in Table 8.
Dataset Source Documents Summary
Multi-News Whether a sign of a good read; or a comment on the ’pulp’ nature of some genres of fiction, the Oxfam second-hand book charts have remained in The Da Vinci Code author’s favour for the past four years. NEWLINE_CHAR NEWLINE_CHAR Dan Brown has topped Oxfam’s ’most donated’ list again, his fourth consecutive year …
A charity shop is urging people to stop donating The Da Vinci Code after becoming overwhelmed with copies. NEWLINE_CHAR NEWLINE_CHAR The Oxfam shop in Swansea has been receiving an average of one copy of the Dan Brown novel a week for months, leaving them with little room for any other books. NEWLINE_CHAR NEWLINE_CHAR Staff who are struggling to sell copies of the book have put a note up in the store saying they would rather donors hand in their vinyl instead.
– The Da Vinci Code has sold so many copies—that would be at least 80 million—that it’s bound to turn up in book donation piles. But at one charity shop in the UK, it’s been donated so heavily that the shop has posted a sign propped up on a tower of Da Vinci Code copies that reads: Ÿou could give us another Da Vinci Code… but we would rather have your vinyl!T̈he manager of the Oxfam shop in Swansea tells the Telegraph that people are laughing and taking pictures of the sizable display: Ï would say that we get one copy of the book every day.Ḧe says people buy them öccasionally,b̈ut with vinyl sales up 25% in the past year, they’d rather take records. Dan Brown’s book isn’t the only one that shops like Oxfam struggle to re-sell. Last year, Oxfam was hit with a large and steady supply of Fifty Shades of Grey, and it similarly begged donors: P̈lease—no more.B̈ut Brown has a particular kind of staying power. The Da Vinci Code was published in 2003, and within six years Brown had booted John Grisham from the No. 1 slot on the list of writers whose books were most often donated to Oxfam’s 700 shops, reported the Guardian at the time. The Independent in 2012 reported Brown’s best-seller was the most-donated book for the fourth year running. (See why Dan Brown took heat from the Philippines.)
@@
Multi-XScience Despite the huge amount of recent research efforts on entity resolution (matching) there has not yet been a comparative evaluation on the relative effectiveness and efficiency of alternate approaches. We therefore present such an evaluation of existing implementations on challenging real-world match tasks. We consider approaches both with and without using machine learning to find suitable parameterization and combination of similarity functions. …
@@
For accurate entity linking and deduplication, early works on the subject examined many methods such as cosine similarity match, distance based match, TF IDF, and Soundex. The well known similarity measures for entity linking are summarized and reviewed in cite ; and the work in cite presents a comparative evaluation of some existing works.
ConvoSumm Title: Hand from tonight. Subreddit: poker Playing at my local pub league. Villian is loose aggressive. Sorry if this is bad formatting just quickly whipped it up for your opinions. |NEWLINE| Button: ( 30000) |NEWLINE| SB: Villian: (50000) …
@@
A lot of commenters debate the semantics and word choice of another comment. Some commenters say that the poster should shove even though other players will not like that move very much. Other commenters discuss the differences of some poker terminology.
Table 3: Examples from each of the three datasets: Multi-News, Multi-XScience, and ConvoSumm-Reddit.
@@
MultiNews Test 5621 2.73 (1.08) 217.00 (67.20) 982.98 (1779.55)
Train 44964 2.76 (1.12) 218.27 (68.04) 982.95 (1851.20)
Val 5621 2.70 (1.04) 216.72 (67.67) 980.06 (1163.94)

### v2

@@ 4 Results
We analyze the results in the framework of our three axes: dataset domain, model type, and evaluation metric. [^7] The domain-transfer analysis is in Appendix D. See Appendix˜E for the detailed per-model and dataset results. To study evaluation metrics, we analyze the performance of the ground-truth dataset. Complete summary examples from each model are in Appendix Figures˜14, 16, 13, 12 and 15.
4.1 MDS Model Generalizability
To study trends specific to each model type, we aggregated the metrics across all the datasets (Table˜1). We also look at correlations between metrics in Appendix Figure˜8. Instead of judging performance by the raw metrics, we look at the relative change in performance from the in-domain setting to the out-of-domain setting (e.g., training on ConvoSumm-Reddit and testing on Multi-XScience). The zero-shot baseline with Llama 3.1 is compared to the in-context Llama 3.1 performance for each dataset. A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data.
@@ 4.1 MDS Model Generalizability
Llama 3.1 outperforms all fine-tuned smaller models in quality and factuality metrics.
Across all datasets in the in-domain setting, Llama generates the most factual and highest-quality summaries according to the automatic metrics. This trend holds during domain transfer, with minimal change in performance (+1+1+ 1 to +4+4+ 4). The exception to the higher performance trend is with the compression metrics and lexical-similarity metrics. Llama has the lowest compression, similar to PRIMERA, as summaries are only $\%5$ compressed.
@@
Surprisingly, Llama has a similar performance at generating summaries with (in-context learning, ICL) or without being shown an example (zero-shot). The Llama-generated summaries change very little when shown examples versus when no examples are provided (Appendix Figures˜15 and 16). This indicates a strong bias to its pre-training versus “learning” from the example. These results could differ with the addition of more examples (at the expense of a smaller context window available for the source documents) or a larger model.
Not all domains benefit from Extract-then-Summarize models.
[…]
@@
With all models, including Llama 3.1 with a 131K context window, there were issues fitting all source documents for every example comfortably in the context windows. As mentioned in Section˜3, PRIMERA “handles” this issue by evenly dividing the context window among the source documents and truncating the ends. The assumption that the beginning of documents is more important than the end does not hold for all domains and has only been identified in News (Dey et al., 2020). While ConvoSumm was able to fit all source documents from ConvoSumm-Reddit, it did have to truncate documents from 25 samples in Multi-XScience and 699 samples from MultiNews. PRIMERA, with a context window double the size of ConvoSumm (2K vs 4K), was able to fit all Conversation and Science samples, but truncated 835 Multi-News examples. Llama only had issues with 3 samples from Multi-News.
The number of source documents does not have strong correlations with other metrics.
Surprisingly, the number of source documents is weakly and inconsistently correlated with factuality and coherence metrics (0.1-0.1- 0.1 to 0.5-0.5- 0.5). This is different than expected, as model performance is expected to decrease with longer context and more information to process.
All models are fluent.
Unsurprisingly, all models have high fluency scores (0.70+0.70+0.70 +) since all models are of size BART-Large and above. While fluency is still an important attribute to check, it might not be as important as it once was.
Word-based similarity metrics are not correlated with Relevancy.
While UniEval-Relevancy measures the similarity between a generated and reference summary (“Is this summary relevant to the reference?”), it does not correlate with the semantic and lexical metrics like MoverScore and ROUGE (0.1-0.1- 0.1 to 0.10.10.1). Whereas MoverScore and ROUGE are highly correlated (0.60.60.6 to 0.80.80.8).
@@ 4.2 Evaluation of Evaluation
To isolate patterns of the evaluation metrics, we look at the automated metric values on the ground-truth reference datasets (Appendix Table˜3) and their correlations (Figure˜3). See Appendix˜E for detailed correlation figures.
Figure 3: Pearson correlations across the ground-truth for all evaluated datasets. The black squares section off intra-correlations within Prometheus and UniEval, and the red square highlights the correlations between them.
UniEval and Prometheus have a weak to moderate correlation.
Over the three datasets, the UniEval and Prometheus metrics have weak to moderate correlations (0.30.30.3 to 0.50.50.5). While Prometheus is $60x$ larger than UniEval (T5-Large), a stronger positive correlation is expected when they are evaluating the same quality (e.g., fluency). The extreme outlier is UniEval-Relevance, which has poor to no correlation with any other metric (0.00.00.0 to 0.20.20.2).
Metric names are not consistent across evaluators and benchmarks.
The definition of “relevance” differs between UniEval and SummEval, “Is this summary relevant to the reference?” versus “How well does the summary capture the key points of the documents?”. This discrepancy explains the 0.00.00.0 correlation between UniEval-Relevance and Prometheus-Relevance. The other metrics have similar definitions and are more correlated, as noted above. See Appendix Figures˜10 and 11 for dimension definitions.
Multi-dimensional evaluators are slightly intra-correlated.
Within Prometheus, the dimensions have slight (0.40.40.4) to moderate (0.70.70.7) correlations with each other, specifically coherence, consistency, and factuality (0.60.60.6 to 0.70.70.7). The UniEval dimensions are less intra-correlated, (0.00.00.0 to 0.40.40.4). This discrepancy in intra-correlation could indicate that Prometheus has difficulty evaluating along a single dimension and instead allows various quality aspects to impact the assigned score. This issue persisted despite adding instructions only to evaluate the desired dimension (see Appendix˜C).
@@
Evaluating for attributes that depend on the source documents (e.g., Consistency) is most affected by the long-document nature of multi-document summarization. The options are either to 1) truncate the documents, which can greatly alter the scores depending on which documents are included, or 2) calculate the metric for every source document and aggregate the scores. More work is needed to determine whether truncating or aggregating per-document scores over- or under-estimates model performance. As for the now-popular “LLM-as-a-Judge” models, like Prometheus, while they have less of a context-window issue, they have a higher computational cost. The reference-based metrics (primarily similarity-based) are not affected by scale because they only compare the ground-truth and generated summaries.
@@ 5 Discussion and Conclusion
In most work that surveys or proposes multi-document summarization (MDS) systems, there is little overlap between the analysis of model performance and the shortcomings of popularly used summarization evaluation metrics. We bridged this gap through experiments and analysis of not only the zero-shot domain transfer ability of MDS systems but also a critical examination of their evaluation methodologies. We analyzed three components of the MDS pipeline for domain transfer: 1) the domain of the training data, 2) the model type, and 3) the common evaluation metrics, including their inter-correlations and the nuances of LLM-as-a-judge approaches.
In conclusion, this work underscores the intricate relationship between MDS model architectures, the stylistic properties of data domains, and the evaluation metrics used to gauge performance. For models, a reduced stylistic bias and the ability to handle extensive context are beneficial for domain transfer. For evaluation, our findings strongly advocate for a multi-faceted approach, exercising caution with lexical metrics for advanced models, and a keen awareness of the specific definitions, potential biases, scaling limitations, and inter-correlation patterns of even sophisticated trained evaluators. Future research should continue to develop more robust and reliable evaluation strategies that are sensitive to the nuances of abstractive summarization across diverse domains and model capabilities.
@@
AI Assistants were used in the preparation of this work, specifically for paper revision, table formatting, and figure editing.
Acknowledgements
@@ Appendix A Dataset Details
We include three MDS datasets: Multi-News+ (Choi et al., 2024), Multi-XScience (Lu et al., 2020)), and ConvoSumm-Reddit (Fabbri et al., 2021a). We keep the original train, validation, and test splits as the original papers for all datasets. See Figures˜6, 4 and 5 for an example from each dataset and Table˜2 for summary statistics.
Multi-News+.
This dataset has 37,057/4,576/4,648 examples for train, validation, and test sets, respectively. We keep the NEWLINE_CHAR symbol in the documents.
@@ Appendix B Training Details
Unless specified otherwise, each model was trained on a single Nvidia A100 GPU with 80GB of memory. The training hyperparameters for PRIMERA, SummN, and ConvoSumm-arg-filtered are in Tables˜7, 6 and 8, respectively. Below are the prompts and details for running Llama for the zero-shot and in-context 1-shot comparisons.
@@
Due to the large size of Llama 8B, we hosted it as a service with vLLM on 4 Nvidia A100 80GB GPUs (Kwon et al., 2023). [^8] https://docs.vllm.ai We followed model examples on how to interact for the summarization task (Table˜5).
Appendix C Evaluation Details
Multi-dimensional analyses with Prometheus.
We developed rubrics for Prometheus based on definitions set forth by prior work. We altered the definitions slightly for the multi-document use case. For all criteria, i.e., description for model feedback, we use the same score rubric except for Factuality and Coverage:
  1. Score 1:
    Very poor
  2. Score 2:
    Poor
  3. Score 3:
    Barely acceptable
  4. Score 4:
    Good
  5. Score 5:
    Very Good
The prompt format is in Figure˜9 and the rubrics are in Figure˜10.
The rubric for Coverage, inspired by Wolhandler et al. (2022b):
  1. Score 1:
    The summary misses most salient information and/or fails to draw from many or all of the source documents.
  2. Score 2:
    The summary includes some salient information, but frequently omits key points or fails to adequately represent all source documents.
  3. Score 3:
    The summary covers a good portion of the salient information and represents most source documents, though some important details or content from some documents may be missing or underrepresented.
  4. Score 4:
    The summary incorporates most salient information from all source documents, with only minor omissions of less critical details, and adequately represents all documents.
  5. Score 5:
    The summary comprehensively and accurately captures all salient information from all source documents, providing a well-rounded representation that integrates content from all sources effectively.
The rubric for Factuality, taken from the Prometheus code repository:
  1. Score 1:
    The model’s responses are mostly incorrect or based on unfounded information.
  2. Score 2:
    The model sometimes provides factually correct responses, but inaccuracies are common.
  3. Score 3:
    The model generally provides factually correct information, though some errors occur.
  4. Score 4:
    The model often provides factually accurate information with only occasional minor errors.
  5. Score 5:
    The model consistently provides responses that are factually correct and well-supported by evidence.
Appendix D MDS Domain Generalizability
We compare the results from training and testing on a single domain to training on a domain and testing on the others (“From”), and training on other domains and testing on the domain (“To”). “Ref” refers to the results when running the metrics on the ground-truth data. See Table˜3.
Models trained on ConvoSumm-Reddit transfer well to other domains.
On average, models trained on ConvoSumm-Reddit and tested on the News and Science datasets see minimal changes in performance across quality and ground-truth similarity metrics. The exception to this trend is the large decrease in MoverScore, which is more sensitive to a performance drop of a hundredth of a decimal due to its often low scores (e.g., 0.050.050.05). [^9] We discuss MoverScore further in Section 4.2.
Transferring from Multi-News+ leads to a decrease in factuality.
Compared to transferring from the other datasets, out-of-domain summaries generated with a model trained on Multi-News+ contain more factual errors (8-11% more). This is most likely due to the large style bias of smaller models, which tend to hallucinate quotes (see Figure˜13 for an example). This is different from the findings of (Belém et al., 2025), who found that models (LLMs) hallucinate more on conversation data than news data.
Most metrics are impacted by models’ bias to the style of the training dataset.
[…]
Appendix E Raw Results
@@
The domain-transfer results in Table˜1 show the summarized relative performance of the models. The detailed performance across domains and models are in Table˜4.
Example from Multi-News+
Source   Documents
Washington ( ap ) — as a crucial second sign-up season gears up, the obama administration said sunday that healthcare.gov is stable and working well, a far cry from last year’s frozen computer screens and frustrated customers. Health and human services secretary sylvia burwell smiles while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,... ( associated press ) health and human services secretary sylvia burwell gestures while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,... ( associated press ) health and human services secretary sylvia m. Burwell said she expects " strong and healthy growth " for 2015. About 7 million people are signed up, and burwell expects to grow that by 2 million more or so. The congressional budget office has projected a total of 13 million enrolled for 2015, and some see the administration as trying to lower expectations. Burwell told nbc’s " meet the press " that 100,000 people had submitted new applications this weekend via the federal website serving 37 states. That’s a big difference from last year, when only a handful of customers managed to enroll on the first day. Burwell also said that a half-million people who already have coverage through the program were able to log into their accounts this time. There were reports saturday that returning customers had problems, but some of that may have been confusion trying to remember user names and passwords. Administration spokesman aaron albright said sunday he had not seen any indication of problems. Burwell said call centers have taken 100,000 calls, another indication of consumer interest. President barack obama noted the improvements. " healthcare.gov works really well now, " he said. Healthcare.gov is an online marketplace that offers subsidized private coverage to people who don’t have health insurance on the job. Because of political opposition and
Getty burwell: 100,000 new obamacare applications secretary of health and human services sylvia mathews burwell said on sunday 100,000 people submitted new applications for obamacare in the first days of the second open enrollment period. Speaking on nbc’s "meet the press, " burwell offered an update to the numbers provided saturday when she said 23,000 applied in the first eight hours of the new open enrollment period. Story continued below in addition to new applicants, 500,000 customers were able to log on to healthcare.gov, and 1 million people have gone "window shopping " to compare insurance plans and prices over the past week. The numbers offer a stark contrast to the first enrollment period a year ago when healthcare.gov launched with severe technical problems that prevented people from applying. Authors:
Summary After last year’s dismal debut healthcare.gov is feeling much better now, and is busily connecting americans with health care plans as the second sign-up period opened yesterday, health and human services secretary sylvia burwell said today. The site took in 100,000 new applications in its first 24 hours, reports politico, saw 500,000 log in to their accounts, and 1 million more went " window-shopping " for plans. Call centers took 100,000 calls. Burwell said she expects " strong and healthy growth " this year — about another 2 million people to add to the 7 million already signed up, per the ap. A white house spokesman said today there was no indication of problems. " healthcare.gov works really well now, " said president obama.
Figure 4: Example from the Multi-News+ dataset.
Example from Multi-XScience
Source Documents
@@
Despite the huge amount of recent research efforts on entity resolution (matching) there has not yet been a comparative evaluation on the relative effectiveness and efficiency of alternate approaches. We therefore present such an evaluation of existing implementations on challenging real-world match tasks. We consider approaches both with and without using machine learning to find suitable parameterization and combination of similarity functions. …
@@
Summary For accurate entity linking and deduplication, early works on the subject examined many methods such as cosine similarity match, distance based match, TF IDF, and Soundex. The well known similarity measures for entity linking are summarized and reviewed in
citekey; and the work in
citekey presents a comparative evaluation of some existing works.
Figure 5: Example from the Multi-XScience dataset.
Example from ConvoSumm-Reddit
Source Documents
Title: Hand from tonight. Subreddit: poker Playing at my local pub league. Villian is loose aggressive. Sorry if this is bad formatting just quickly whipped it up for your opinions. Button: ( 30000) SB: Villian: (50000) …
@@
Summary A lot of commenters debate the semantics and word choice of another comment. Some commenters say that the poster should shove even though other players will not like that move very much. Other commenters discuss the differences of some poker terminology.
Figure 6: Example from the ConvoSumm-Reddit dataset.
@@
MultiNews+ Test 4648 2.53 (0.88) 230.94 (69.57) 226.07 (64.15)
Train 37057 2.55 (0.91) 233.54 (69.98) 224.30 (64.96)
Val 4576 2.51 (0.86) 231.45 (69.64) 226.59 (63.35)